# Towards Pervasive Distributed Agentic Generative AI - A State of The Art

### v1

@@ 4.2. LLMs Alignment Strategies
However, a significant drawback of full fine-tuning is that each downstream task necessitates learning a new set of parameters, $\Phi$, whose size is equivalent to the original model’s parameter set, $\Phi_{0}$. In resource-limited scenarios, such as pervasive computing environments, this approach becomes prohibitively complex and computationally expensive. To mitigate these challenges, Parameter-Efficient Fine-Tuning (PEFT) methods have emerged. These strategies are specifically designed to enable model adaptation even under stringent resource constraints. Among prominent PEFT techniques are Low-Rank Adaptation (Hu et al., 2022) (LoRA) and its quantized variant, QLoRA (Dettmers et al., 2023). These methods enable efficient fine-tuning by learning a significantly smaller set of task-specific parameters, denoted as $\Theta$ , where $|\Theta|\ll|\Phi_{0}|$. Consequently, the task of determining the weight update $\Delta\Phi$ is reframed as an optimization problem over the smaller parameter set $\Theta$:
@@
This approach significantly reduces the memory and computational overhead associated with fine-tuning, making model adaptation more feasible in resource-constrained settings. For instance InfiGUI Agent (Liu et al., 2025) uses Full Supervised Fine-Tuning in two stages. In Stage 1, they collected diverse page contents and vision-language datasets, using screen coordinates, to train basic skills like page contents understanding and grounding. In Stage 2, they synthesized new data to teach two advanced reasoning skills: hierarchical reasoning (strategic + tactical task planning) and expectation-reflection reasoning (self-correction from outcomes). For dataset creation, they used both real page contents datasets and synthetic SFT data generated from multimodal LLMs. They focused on ”Reference-Augmented Annotations” to precisely link visual elements and textual reasoning. ReachAgent (Wu et al., [n. d.]) uses a first stage of Supervised Fine-Tuning (SFT). They build three datasets: Page Navigation, Page Reaching, and Page Operation. Each dataset focuses on different subtasks such as reaching pages, operating within pages, and navigating multi-step tasks. They generated tasks and step-by-step labels using various LLMs. The SFT trains the model to understand full page contents flows and improve multi-step planning before any reinforcement learning. This stage ensures the agent can solve subtasks accurately before optimizing for full-task preferences in later RL training. MaViLa’s Supervised Fine-Tuning (SFT) to perform visual scene understanding, anomaly detection, and manufacturing reasoning using using LoRA on a Vicuna 13B model (Fan et al., 2025). For dataset creation, they collected real-world and schematic manufacturing images, each manually captioned. They generated instruction-response pairs by prompting GPT-4, distinguishing between general and domain-specific questions. For domain-specific instructions, they used Retrieval-Augmented Generation (RAG) to ground answers in manufacturing knowledge. Instructions were classified by complexity, reasoning need, and domain specificity to ensure high-quality fine-tuning data. Finally in (Yonekura et al., 2024) researchers used Supervised Fine-Tuning (SFT) to customise LLMs for stable activity generation in smart home simulations. To create the fine-tuning dataset, they collected labeled examples of human-like daily schedules and activity outputs. Fine-tuning focused on ensuring structured outputs (like JSON) to prevent simulator crashes from unstructured LLM replies. The SFT enhanced the model’s ability to generate realistic, context-aware daily activities for virtual smart home agents. As a result, they achieved a 4.3% improvement in simulation stability and efficiency compared to baseline prompting.
@@
Another effective approach for aligning LLMs through Fine-Tuning is Direct Preference Optimization (DPO) (Rafailov et al., 2023). This method offers a more efficient approach to instruction-tuning. It directly trains the LLM agent on pairs of preferred (”winner”) and less-preferred (”loser”) responses to an instruction. Critically, DPO achieves this by fine-tuning LLMs without the need for an explicit reward model, relying solely on these direct preference comparisons between output pairs. The DPO algorithm begins by sampling completions $y_{1},y_{2}\sim\pi_{\text{ref}}(\cdot\mid x)$ for each prompt $x$, and then labels these completions based on human preferences to construct an offline preference dataset:
@@
For instance ReachAgent (Wu et al., [n. d.]) uses Direct Preference Optimization (DPO) to refine its decision-making in mobile GUI tasks. DPO is used without needing explicit numeric rewards, instead relying on preference pairs indicating which actions are better. These preferences are constructed using a 4-level reward ranking (Golden ¿ Longer ¿ Incomplete ¿ Invalid) based on how effectively and efficiently page contents flows complete the task. During training, the model learns to prefer actions that lead to more optimal flows—those that are both task-completing and concise. This preference data is fed into the DPO loss function, which guides the policy ($\pi_{\theta}$) to align closer to preferred behaviors while softly deviating from the supervised fine-tuned policy ($\pi_{SFT}$). This enhances the model’s ability to generate efficient and successful page contents flows without needing exact matches to gold actions.
@@
Federated Learning (FL) is a machine learning paradigm that enables the training of a unified model across numerous decentralized edge devices or servers, each holding local data, without the need to exchange these sensitive datasets (Bimpas et al., 2024). Essentially, FL allows for collaborative algorithm training on distributed data sources while preserving data locality. One of the core principle of FL is to empower devices, including smartphones and IoT devices, to collectively learn a shared predictive model, distributing the training computation between multiple devices. For instance in FedMobileAgent framework (Wang et al., 2025b), federated learning was used to collaboratively train mobile agents across decentralized user data while preserving privacy. First each user locally collects data $D_{k}=\left\{\langle T,a_{i},s_{i}\rangle\right\}_{i=1}^{n}$ via Auto-Annotation, where $T$ is a task instruction, $a_{i}$ an action, and $s_{i}$ a screenshot. Then local training updates the model using stochastic gradient descent:
$M_{k}^{(l,r+1)}=M_{k}^{(l,r)}-\eta\nabla\ell\left(M_{k}^{(l,\tau_{k})};T,s,a\right)\hskip 5.0pt\text{where }M_{k}^{(l,\tau_{k})}\text{ is sent to the server, }M^{(l+1)}=\sum_{k\in S^{l}}\omega_{k}M_{k}^{(l)},\text{ and }\omega_{k}=\frac{n_{k}^{*}}{\sum_{k\in S^{l}}n_{k}^{*}}.$
@@ 4.3.1. Memory
This memory saves the intermediate steps within its current execution cycle. These ongoing steps are readily accessible at each decision-making stage, allowing the agent to reason about the next action in a context-aware manner. This mechanism significantly reduces the likelihood of the agent becoming trapped in repetitive states or endlessly cycling through the same sequence of actions. For instance in (Wang et al., 2024d) the agent employs short-term memory to meticulously record all actions performed and the corresponding state changes as it interacts with the environment. This dynamically observed information then directly informs the next operational decisions. Also in (Li et al., 2024b) the agent uses short-term memory to maintain a log of the current session’s context, including all prior reasoning steps, executed actions, and received observations. This memory is structured and iteratively updated after each tool execution, ensuring a coherent and up-to-date understanding of the ongoing situation. This allows the LLM to generate contextually relevant thoughts and actions at every stage of its processing pipeline. Crucially, this short-term memory is session-bound and discarded upon the completion of the inference process, guaranteeing real-time decision traceability without necessitating long-term storage of transient data.
@@
Once an agent concludes its execution, regardless of whether the final goal was achieved, all the intermediate steps and experiences accumulated during the session are consolidated. Through various summarization or simplification techniques, this information is then archived in Long-Term Memory for future reference and learning. For instance in (Rivkin et al., 2024) the agent leverages long-term memory to store a comprehensive history of user interactions with the agent. It employs a vector database, using the MiniLM embedding model for efficient storage and retrieval of semantically similar past executions when confronted with new tasks. Additionally, high-level summaries of user-agent interactions are stored to build a dynamic and holistic understanding of individual user preferences over time. Mobile-Agent-E (Wang et al., 2025a) manages memory through a persistent long-term storage, focusing on two key elements derived from past experiences: Tips and Shortcuts. Tips represent generalizable insights gleaned from prior tasks, providing guidance for both high-level strategic planning and low-level action execution. Shortcuts are reusable sequences of actions identified for frequently occurring subroutines. Following each completed task, two ”Experience Reflectors” analyse the entire interaction history to update existing Tips and Shortcuts or generate new ones. These refined or novel insights are then using by the ”Manager” for future planning and the ”Operator” for the next action execution. Finally in (Lee et al., 2024), the agent employs a hierarchical memory system that stores tasks as ordered sequences of subtasks and actions, directly linked to specific application screens. Each screen is represented as a node containing a set of available subtasks, with edges denoting transitions between subtasks triggered by related low-level actions. Tasks are saved in a parameterized function-call format, enabling flexible reuse across different contexts. During execution, the system can recall previously encountered tasks or subtasks and adapt them to new situations through attribute matching and in-context learning techniques. The memory is dynamically updated based on user feedback or the agent’s own self-correction mechanisms. As previously discussed, long-term memory serves as a repository of both successful and failed executions, enabling the agent to learn and improve over time. Two core strategies are commonly employed for populating this valuable resource:
@@
  • Exploratory Phase: Often, before being deployed for user interaction, an agent goes through a dedicated exploratory phase. During this stage, the agent is encouraged to actively explore its environment (e.g., the functionalities of various applications on a device). The primary objective is to develop a foundational understanding of the interaction mechanisms, identify the available tools and their functionalities, and learn the effects of different actions and strategies for overcoming specific challenges. This proactive exploration ensures that, when the agent is eventually used for real tasks, it already possesses a significant knowledge to more reliably navigate the environment and find the correct path to achieve its goals. For instance MobileGPT (Lee et al., 2024) analyzes app screens offline, extracting UI layout and simplifying it to HTML. The LLM identifies subtasks, formats them as function calls, and caches them for efficient live execution. AutoDroid (Wen et al., 2024a) explores apps by random UI interactions, building a UI Transition Graph summarizing states and elements as simulated tasks, stored in memory. Also AutoDroid-V2 (Wen et al., 2024b) constructs structured documents from page contents traces, abstracting states and transitions.
@@
  • Test Phase: In this scenario, the agent continuously learns from its interactions with the user during live testing. Both successful and failed execution attempts are stored in long-term memory. This continuous learning process allows the agent to gradually refine its strategies and improve its performance as the user engages in more tasks. For instance IDS Agent (Li et al., 2024b) uses long-term memory to resolve ambiguous situations by retrieving past sessions similar to the current one, incorporating their reasoning and outcomes for better decisions. Also the LiMeDa Framework (Chen et al., 2024b) stores summarized data from completed vehicle tasks (route, time, energy) in memory, managing it to prevent overload. Upon receiving a new task, LiMeDa retrieves relevant past experiences to improve decision efficiency and avoid repeating errors.
@@ 4.3.2. Reasoning and Planning
Following the perception of the system state and the memory module’s contents, including previous actions or complete past executions, the agent proceeds through Reasoning and Planning. The Reasoning step bring out a logical thinking process from the LLM, facilitating the formulation of an effective plan. For instance Mobile Agent-E (Wang et al., 2025a) implements an Action-Reflection module that analyzes the state before and after an action, along with the action itself, to determine if the outcome aligns with the expected goal. This module categorizes action outcomes into: Successful or partially successful (outcome matches expectation), Failed because result leads to an incorrect state, and Failed because an action produces no observable change. Also InfiGUI Agent (Liu et al., 2025) operates in a three-step cycle: Reasoning (performing hierarchical reasoning), Action generation (writing the next action and its anticipated outcomes), and Reflection (analyzing the resulting state to evaluate if the expected results were achieved and generating a textual summary of this reflection). A popular reasoning technique is Chain of Thought (CoT) prompting (Wei et al., 2022) , that allows the agent to tackle complex reasoning tasks by breaking them down into a sequence of explicit steps, often initiated by prompts such as ”Let’s think step by step”. For instance Mavila (Fan et al., 2025) , used for process automation, employs CoT reasoning to decompose tasks into a series of sequential steps and in the SAGE framework (Rivkin et al., 2024), planning (the decomposition of a high-level goal into substeps) is managed using CoT reasoning in conjunction with the ReAct (Yao et al., 2023) design pattern. Tool instructions and formatting guidelines encourage the LLM to first outline a plan before specifying the execution details. Also AutoDroid (Wen et al., 2024a) fine-tunes a small language model (SLM) to reason with a zero-shot CoT approach, using a structured format.
@@
The other step is Planning, where, informed by prior events and the reasoning process, the agent decides on the next action(s) to take. For instance in the IDS Agent (Li et al., 2024b), planning involves selecting the appropriate tool based on the information in memory and the analyzed traffic data. The agent then decides, based on classification results, whether to trigger an alert, thus identifying a potential security threat. Mobile Agent-E (Wang et al., 2025a) features a Manager module responsible for generating a plan broken down into subtasks. At each step, the Manager considers the initial user query, the current screenshot, the previous overall plan, the preceding subgoal, the current progress status, available Shortcuts from long-term memory, and any relevant notes. It then updates the overall plan and identifies the next immediate subgoal to pursue. Finally in (Wang et al., 2024d) the agent employs three distinct agent roles for mobile device interaction. The Planning Agent determines the task progression using historical data. The Decision Agent observes relevant content from past screens via memory to generate actions and update the memory. The Reflection Agent evaluates if actions meet expectations by comparing screen states and initiates corrective re-execution if needed.

### v2

@@ 5.2. LLMs Alignment Strategies
However, a significant drawback of full fine-tuning is that each downstream task necessitates learning a new set of parameters, $\Phi$, whose size is equivalent to the original model’s parameter set, $\Phi_{0}$. In resource-limited scenarios, such as pervasive computing environments, this approach becomes prohibitively complex and computationally expensive. To mitigate these challenges, Parameter-Efficient Fine-Tuning (PEFT) methods have emerged. These strategies are specifically designed to enable model adaptation even under stringent resource constraints. Among prominent PEFT techniques are Low-Rank Adaptation (hu2022lora) (LoRA) and its quantized variant, QLoRA (dettmers2023qlora). These methods enable efficient fine-tuning by learning a significantly smaller set of task-specific parameters, denoted as $\Theta$ , where $|\Theta|\ll|\Phi_{0}|$. Consequently, the task of determining the weight update $\Delta\Phi$ is reframed as an optimization problem over the smaller parameter set $\Theta$:
@@
This approach significantly reduces the memory and computational overhead associated with fine-tuning, making model adaptation more feasible in resource-constrained settings. For instance InfiGUI Agent (liu2025infiguiagent) uses Full Supervised Fine-Tuning in two stages. In Stage 1, they collected diverse page contents and vision-language datasets, using screen coordinates, to train basic skills like page contents understanding and grounding. In Stage 2, they synthesized new data to teach two advanced reasoning skills: hierarchical reasoning (strategic + tactical task planning) and expectation-reflection reasoning (self-correction from outcomes). For dataset creation, they used both real page contents datasets and synthetic SFT data generated from multimodal LLMs. They focused on ”Reference-Augmented Annotations” to precisely link visual elements and textual reasoning. ReachAgent (wureachagent) uses a first stage of Supervised Fine-Tuning (SFT). They build three datasets: Page Navigation, Page Reaching, and Page Operation. Each dataset focuses on different subtasks such as reaching pages, operating within pages, and navigating multi-step tasks. They generated tasks and step-by-step labels using various LLMs. The SFT trains the model to understand full page contents flows and improve multi-step planning before any reinforcement learning. This stage ensures the agent can solve subtasks accurately before optimizing for full-task preferences in later RL training. MaViLa’s Supervised Fine-Tuning (SFT) to perform visual scene understanding, anomaly detection, and manufacturing reasoning using using LoRA on a Vicuna 13B model (fan2025mavila). For dataset creation, they collected real-world and schematic manufacturing images, each manually captioned. They generated instruction-response pairs by prompting GPT-4, distinguishing between general and domain-specific questions. For domain-specific instructions, they used Retrieval-Augmented Generation (RAG) to ground answers in manufacturing knowledge. Instructions were classified by complexity, reasoning need, and domain specificity to ensure high-quality fine-tuning data. Finally in (yonekura2024generating) researchers used Supervised Fine-Tuning (SFT) to customise LLMs for stable activity generation in smart home simulations. To create the fine-tuning dataset, they collected labeled examples of human-like daily schedules and activity outputs. Fine-tuning focused on ensuring structured outputs (like JSON) to prevent simulator crashes from unstructured LLM replies. The SFT enhanced the model’s ability to generate realistic, context-aware daily activities for virtual smart home agents. As a result, they achieved a 4.3% improvement in simulation stability and efficiency compared to baseline prompting.
@@
Another effective approach for aligning LLMs through Fine-Tuning is Direct Preference Optimization (DPO) (rafailov2023direct). This method offers a more efficient approach to instruction-tuning. It directly trains the LLM agent on pairs of preferred (”winner”) and less-preferred (”loser”) responses to an instruction. Critically, DPO achieves this by fine-tuning LLMs without the need for an explicit reward model, relying solely on these direct preference comparisons between output pairs. The DPO algorithm begins by sampling completions $y_{1},y_{2}\sim\pi_{\text{ref}}(\cdot\mid x)$ for each prompt $x$, and then labels these completions based on human preferences to construct an offline preference dataset:
@@
For instance ReachAgent (wureachagent) uses Direct Preference Optimization (DPO) to refine its decision-making in mobile GUI tasks. DPO is used without needing explicit numeric rewards, instead relying on preference pairs indicating which actions are better. These preferences are constructed using a 4-level reward ranking (Golden ¿ Longer ¿ Incomplete ¿ Invalid) based on how effectively and efficiently page contents flows complete the task. During training, the model learns to prefer actions that lead to more optimal flows—those that are both task-completing and concise. This preference data is fed into the DPO loss function, which guides the policy ($\pi_{\theta}$) to align closer to preferred behaviors while softly deviating from the supervised fine-tuned policy ($\pi_{SFT}$). This enhances the model’s ability to generate efficient and successful page contents flows without needing exact matches to gold actions.
@@
Federated Learning (FL) is a machine learning paradigm that enables the training of a unified model across numerous decentralized edge devices or servers, each holding local data, without the need to exchange these sensitive datasets (bimpas2024leveraging). Essentially, FL allows for collaborative algorithm training on distributed data sources while preserving data locality. One of the core principle of FL is to empower devices, including smartphones and IoT devices, to collectively learn a shared predictive model, distributing the training computation between multiple devices. For instance in FedMobileAgent framework (wang2025fedmobileagent), federated learning was used to collaboratively train mobile agents across decentralized user data while preserving privacy. First each user locally collects data $D_{k}=\left\{\langle T,a_{i},s_{i}\rangle\right\}_{i=1}^{n}$ via Auto-Annotation, where $T$ is a task instruction, $a_{i}$ an action, and $s_{i}$ a screenshot. Then local training updates the model using stochastic gradient descent:
[…]
@@ 5.3.1. Memory
This memory saves the intermediate steps within its current execution cycle. These ongoing steps are readily accessible at each decision-making stage, allowing the agent to reason about the next action in a context-aware manner. This mechanism significantly reduces the likelihood of the agent becoming trapped in repetitive states or endlessly cycling through the same sequence of actions. For instance in (wang2024mobile) the agent employs short-term memory to meticulously record all actions performed and the corresponding state changes as it interacts with the environment. This dynamically observed information then directly informs the next operational decisions. Also in (li2024ids) the agent uses short-term memory to maintain a log of the current session’s context, including all prior reasoning steps, executed actions, and received observations. This memory is structured and iteratively updated after each tool execution, ensuring a coherent and up-to-date understanding of the ongoing situation. This allows the LLM to generate contextually relevant thoughts and actions at every stage of its processing pipeline. Crucially, this short-term memory is session-bound and discarded upon the completion of the inference process, guaranteeing real-time decision traceability without necessitating long-term storage of transient data.
@@
Once an agent concludes its execution, regardless of whether the final goal was achieved, all the intermediate steps and experiences accumulated during the session are consolidated. Through various summarization or simplification techniques, this information is then archived in Long-Term Memory for future reference and learning. For instance in (rivkin2024aiot) the agent leverages long-term memory to store a comprehensive history of user interactions with the agent. It employs a vector database, using the MiniLM embedding model for efficient storage and retrieval of semantically similar past executions when confronted with new tasks. Additionally, high-level summaries of user-agent interactions are stored to build a dynamic and holistic understanding of individual user preferences over time. Mobile-Agent-E (wang2025mobile) manages memory through a persistent long-term storage, focusing on two key elements derived from past experiences: Tips and Shortcuts. Tips represent generalizable insights gleaned from prior tasks, providing guidance for both high-level strategic planning and low-level action execution. Shortcuts are reusable sequences of actions identified for frequently occurring subroutines. Following each completed task, two ”Experience Reflectors” analyse the entire interaction history to update existing Tips and Shortcuts or generate new ones. These refined or novel insights are then using by the ”Manager” for future planning and the ”Operator” for the next action execution. Finally in (lee2024mobilegpt), the agent employs a hierarchical memory system that stores tasks as ordered sequences of subtasks and actions, directly linked to specific application screens. Each screen is represented as a node containing a set of available subtasks, with edges denoting transitions between subtasks triggered by related low-level actions. Tasks are saved in a parameterized function-call format, enabling flexible reuse across different contexts. During execution, the system can recall previously encountered tasks or subtasks and adapt them to new situations through attribute matching and in-context learning techniques. The memory is dynamically updated based on user feedback or the agent’s own self-correction mechanisms.
As previously discussed, long-term memory serves as a repository of both successful and failed executions, enabling the agent to learn and improve over time. Two core strategies are commonly employed for populating this valuable resource:
@@
  • Exploratory Phase: Often, before being deployed for user interaction, an agent goes through a dedicated exploratory phase. During this stage, the agent is encouraged to actively explore its environment (e.g., the functionalities of various applications on a device). The primary objective is to develop a foundational understanding of the interaction mechanisms, identify the available tools and their functionalities, and learn the effects of different actions and strategies for overcoming specific challenges. This proactive exploration ensures that, when the agent is eventually used for real tasks, it already possesses a significant knowledge to more reliably navigate the environment and find the correct path to achieve its goals. For instance MobileGPT (lee2024mobilegpt) analyzes app screens offline, extracting UI layout and simplifying it to HTML. The LLM identifies subtasks, formats them as function calls, and caches them for efficient live execution. AutoDroid (wen2024autodroidv1) explores apps by random UI interactions, building a UI Transition Graph summarizing states and elements as simulated tasks, stored in memory. Also AutoDroid-V2 (wen2024autodroid) constructs structured documents from page contents traces, abstracting states and transitions.
@@
  • Test Phase: In this scenario, the agent continuously learns from its interactions with the user during live testing. Both successful and failed execution attempts are stored in long-term memory. This continuous learning process allows the agent to gradually refine its strategies and improve its performance as the user engages in more tasks. For instance IDS Agent (li2024ids) uses long-term memory to resolve ambiguous situations by retrieving past sessions similar to the current one, incorporating their reasoning and outcomes for better decisions. Also the LiMeDa Framework (chen2024llm) stores summarized data from completed vehicle tasks (route, time, energy) in memory, managing it to prevent overload. Upon receiving a new task, LiMeDa retrieves relevant past experiences to improve decision efficiency and avoid repeating errors.
@@ 5.3.2. Reasoning and Planning
Following the perception of the system state and the memory module’s contents, including previous actions or complete past executions, the agent proceeds through Reasoning and Planning. The Reasoning step bring out a logical thinking process from the LLM, facilitating the formulation of an effective plan. For instance Mobile Agent-E (wang2025mobile) implements an Action-Reflection module that analyzes the state before and after an action, along with the action itself, to determine if the outcome aligns with the expected goal. This module categorizes action outcomes into: Successful or partially successful (outcome matches expectation), Failed because result leads to an incorrect state, and Failed because an action produces no observable change. Also InfiGUI Agent (liu2025infiguiagent) operates in a three-step cycle: Reasoning (performing hierarchical reasoning), Action generation (writing the next action and its anticipated outcomes), and Reflection (analyzing the resulting state to evaluate if the expected results were achieved and generating a textual summary of this reflection). A popular reasoning technique is Chain of Thought (CoT) prompting (wei2022chain) , that allows the agent to tackle complex reasoning tasks by breaking them down into a sequence of explicit steps, often initiated by prompts such as ”Let’s think step by step”. For instance Mavila (fan2025mavila) , used for process automation, employs CoT reasoning to decompose tasks into a series of sequential steps and in the SAGE framework (rivkin2024aiot), planning (the decomposition of a high-level goal into substeps) is managed using CoT reasoning in conjunction with the ReAct (yao2023react) design pattern. Tool instructions and formatting guidelines encourage the LLM to first outline a plan before specifying the execution details. Also AutoDroid (wen2024autodroidv1) fine-tunes a small language model (SLM) to reason with a zero-shot CoT approach, using a structured format.
@@
The other step is Planning, where, informed by prior events and the reasoning process, the agent decides on the next action(s) to take. For instance in the IDS Agent (li2024ids), planning involves selecting the appropriate tool based on the information in memory and the analyzed traffic data. The agent then decides, based on classification results, whether to trigger an alert, thus identifying a potential security threat. Mobile Agent-E (wang2025mobile) features a Manager module responsible for generating a plan broken down into subtasks. At each step, the Manager considers the initial user query, the current screenshot, the previous overall plan, the preceding subgoal, the current progress status, available Shortcuts from long-term memory, and any relevant notes. It then updates the overall plan and identifies the next immediate subgoal to pursue. Finally in (wang2024mobile) the agent employs three distinct agent roles for mobile device interaction. The Planning Agent determines the task progression using historical data. The Decision Agent observes relevant content from past screens via memory to generate actions and update the memory. The Reflection Agent evaluates if actions meet expectations by comparing screen states and initiates corrective re-execution if needed.